OpenStudy (anonymous):

how would I solve 4x=-12y+4?

OpenStudy (anonymous):

What do you mean by solve?

OpenStudy (anonymous):

you can only solve one variable in term of the other

OpenStudy (anonymous):

solve to graph a linear equation

OpenStudy (anonymous):

then you solve for y.

OpenStudy (anonymous):

12y=4-4x y=(1/3)-(1/3)x

OpenStudy (anonymous):

graph that and you got your answer

OpenStudy (anonymous):

I dont get how you got to that part though

OpenStudy (anonymous):

subtract 4x on both side, add 12y then you divide by 12. 4/12 = 1/3

OpenStudy (anonymous):

that is pretty easy take x zero first y would become 1/3 and now then take y as 0 x would become 1 repeat this for at least 4 values for both x and y you would get points and plot them that would be your graph

OpenStudy (anonymous):

ok I see thank you.

OpenStudy (anonymous):

np